# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### 13 Dec 2023 v1

@@ Abstract
Evolutionary multitasking (EMT) has been attracting much attention over the past years. It aims to handle multiple optimization tasks simultaneously within limited computing resources assisted by inter-task knowledge transfer techniques. Numerous multitask evolutionary algorithms (MTEAs) for solving multitask optimization (MTO) problems have been proposed in the EMT field, but there lacks a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To address this issue, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. It incorporates more than 30 MTEAs, more than 150 MTO problem cases with real-world applications, and more than 10 performance metrics. Moreover, for comparing MTEAs with traditional evolutionary algorithms, we modified more than 30 popular single-task evolutionary algorithms to be able to solve MTO problems in MToP. MToP is a user-friendly tool with a graphical user interface that makes it easy to analyze results, export data, and plot schematics. More importantly, MToP is extensible, allowing users to develop new algorithms and define new problems. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
@@ I Introduction
Evolutionary computation (EC) is inspired by natural evolution and has grown rapidly due to effectiveness and efficiency. With its robustness and ease of use, EC algorithms, also known as evolutionary algorithms (EAs), have achieved great success in solving black-box optimization problems. Designing specific EAs for different complex black-box optimization problems, such as constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6], has also been studied extensively. In recent years, with the increase in computational demands and the development of cloud computing, the use of EAs to handle a large number of optimization tasks simultaneously has gained much focus. This requirement is called evolutionary multitasking (EMT) [7]. A multitask optimization (MTO) problem of EMT with $K$ optimization tasks aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks. It can be formulated as follows:
@@
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. For multi-objective optimization task, its $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$ task. EMT has been successfully applied to engineering scheduling [5], feature selection in machine learning [8], anomaly detection [9], point cloud registration [10], and reinforcement learning [11].
@@
In order to solve multiple optimization tasks faster and better concurrently, researchers have attempted to exploit the similarity among tasks to empower EAs with inter-task knowledge extraction and transfer [12]. Through knowledge transfer, EAs are then able to fully utilize implicit parallelism to obtain better solutions for multiple tasks with fewer computational resources [13]. The first attempt for EMT is the multifactorial EA [14], which proposes an implicit knowledge representation through random mating among optimization tasks. After that, a large number of EAs for solving MTO problems have been proposed. These developed multitask EAs (MTEAs) employ either a multifactorial framework [15, 16] that utilizes a single population for multiple tasks or a multi-population framework [17, 18] that assign separate populations for multiple tasks. In addition, for decision space mapping among different tasks, various techniques in MTEAs such as unified search space [14, 16], autoencoding [19, 20], affine transformation [21, 22], and adversarial generative model [23] have been successively proposed. Due to the significant influence of the amount of knowledge transfer for solving MTO problems, MTEAs with knowledge transfer adaptive control strategies, such as similarity judgment [15, 24], knowledge selection [18, 25], and historical feedback [26, 27], have also been studied. There is a growing number of MTEAs proposed by researchers, yet there is no standardized programming language, code pattern, or software platform for the EMT source codes. This presents difficulties for beginners to enter the field of EMT and for researchers to make convenient experimental comparisons of the algorithms.
As EMT is booming, it has become urgent to provide a convenient and user-friendly comparison software platform for benchmarking the performance of MTEAs. Moreover, the available source code and platform are essential for exploring real-world applications of EMT. Open-source and user-friendly software platforms can greatly facilitate the development of a research field. As for EC, some popular and successful software platforms helped researchers advance the development of EA and evolutionary optimization. PlatEMO [28] is developed for evolutionary multi-objective optimization and benchmarking multi-objective EAs. It facilitates researchers to design multi-objective EAs to deal with various types of multi-objective optimization problems. EvoX [29] is a distributed GPU-accelerated library that helps researchers design parallel EAs to accelerate the solving of complex optimization problems and training for reinforcement learning. However, different from the traditional single-task EC field, solving MTO problems with MTEAs requires simultaneous evolution of multiple optimization tasks with inter-task solution space mapping and knowledge transfer. Additionally, the performance metrics of EMT are diverse, with both single-task metrics and multitask metrics. These special requirements are hard to implement in existing tools and platforms. Moreover, existing research work on EMT presents various code implementations of algorithms and problems. Particularly, EMT has been extended to the subfields of many-task optimization [30, 31], multi-objective multitask optimization [32, 33], competitive multitask optimization [26, 34], and constrained multi-task optimization [35]. The researchers cannot easily benchmark MTEAs on various types of MTO problems.
@@
In order to address these issues, we introduce an open-source MATLAB platform, named MTO-Platform (MToP), for the development of the EMT field. The main contributions of MToP are summarized as follows:
@@
    MToP has an easy-to-use graphical user interface (GUI) including test, experiment, and data-process modules. These modules can assist researchers in realizing problem characteristics, performing comparative experiments, parallel execution, statistically analyzing results, plotting result figures, and processing experimental data. Moreover, MToP provides a modular implementation of algorithms, problems and metrics. On top of that, extensive public application programming interface (API) functions are provided, such as population initialization, function evaluation, evolutionary operator, and environmental selection.
@@
    An abundance of algorithms, problems, and metrics are implemented with public API and included in MToP. More than 30 MTEAs (see Table I) cover single-objective, multi-objective, constrained, many-task, and competitive EMT types. Additionally, to facilitate the comparison of MTEAs with popular traditional EAs, more than 30 single-task EAs (see Table II) of various types, such as genetic algorithm, differential evolution, swarm intelligence, and evolution strategy, are also modified and included in MToP. For benchmark test problems, more than 150 benchmark MTO problems and several real-world applications (see Table III) are incorporated. Finally, various types of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, are also implemented in MToP.
@@
    MToP is extensible to easily adding new code of algorithms, problems, and metrics. After following the coding patterns and implementing functionality based on the public API, new code can be recognized and performed. Since MToP is completely open-source, researchers can utilize existing algorithms, problems, and metrics as the basis for new ideas. Through the open-source website GitHub, MToP will be continuously updated and receive new code [^1] MToP has been continuously updated since Sep. 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
TABLE I: Multitask evolutionary algorithms included in MToP (current version).
MTEA-AD [18] 2022 TEVC Single Multi Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
MTES-KG [46] 2023 TEVC Single Multi Multi-population Yes Multitask ES with knowledge-guided external sampling
MTEA-SaO [48] 2023 INS Single/Multi Multi Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
MO-MFEA [32] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominanted sort and crowding distance
EMT-PD [33] 2021 SMCA Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
TABLE II: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
C2oDE [72] 2019 SMCA Single Constraint Handling Yes Composite DE for constrained optimization
DeCODE [78] 2021 SMCA Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
CEDE-DRL [80] 2023 SWEVO Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
TABLE III: Multitask optimization problems included in MToP (current version).
Problem Case Task Dimension Objective Special Description
SOPM 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
MO-OPF 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
MO-SCP 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
TABLE IV: Metrics included in MToP (current version).
Metric Objective Description
Obj Single Objective value for each task
Obj (MTS) Single Multitask score of Obj for all tasks
Obj (AV) Single Average Obj for all tasks
Obj (CMT) Single Competitive multitask Obj for all tasks
Obj (NBR) Single Number of best Obj result for all tasks
CV Single Constraint violation for each task
FR Single Feasible rate for each task
HV Multi Hypervolume for each task
HV (MTS) Multi Multitask score of HV for all tasks
HV (CMT) Multi Competitive multitask HV for all tasks
IGD Multi Inverted generational distance for each task
IGD+ Multi Improved IGD plus for each task
IGD+ (MTS) Multi Multitask score of IGD+ for all tasks
IGD+ (CMT) Multi Competitive multitask IGD+ for all tasks
Run Time All Algorithm running time for all tasks
The rest of this paper is organized as follows. Section II elaborates on the architecture of MToP. Section III describes approaches for performing experiments with MToP. Section IV introduces how to extend contents of MToP. Finally, the conclusion and future work are given in Section V.
@@ II-A1 Test Module
(a) Function landscapes
(b) Feasible regions
(c) Pareto front
(d) Metric convergence
@@
The test module is designed to facilitate researchers in analyzing the characteristics of MTO problems and algorithms. The test module of MToP provides functions to run the selected algorithm on the selected problem. It provides visualization functions for both single-objective and multi-objective MTO problems. For single-objective MTO problems, the landscapes of the objective function with different tasks in the one-dimensional unified search space can be displayed as in Fig. 1 (a). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be shown as in Fig. 1 (b). For multi-objective MTO problems, the Pareto front figures can be displayed for multiple tasks as in Fig. 1 (c). The metric convergence behavior can be drawn after running algorithms on problems as in Fig. 1 (d).
@@
The workflow of API calls for the execution of an algorithm on a problem is displayed in Fig. 2. The workflow starts with the Start_Button event in MTO_GUI.m, which calls the parameter setting API Set_Parameter in the selected algorithm MFEA.m. Following this, it enters the Run function of the algorithm, which at the beginning calls the Initialization API of the algorithm base class Algorithm.m and later loops through the function evaluation API Evaluation. The Initialization gets the parameter settings of the problem CEC17_MTP1.m to be solved via the problem base class Problem.m. Similarly, the Evaluation gets the specific optimization task object through the problem base class. After the algorithm has run, GUI then displays the resultant performance metrics after calling the metrics function Obj.m via Call_Metric API. All functions, including the subsequent experiment module, in MToP that execute algorithms are based on this procedure. Both algorithm and problem parameters can be modified directly through the GUI of MToP.
@@ II-A2 Experiment Module
TABLE V: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
@@
The experiment module provides functions for performing comprehensive experiments and analyzing results. Users can execute multiple independent runs of multiple algorithms to solve multiple MTO problems. Data such as experimental settings, objective values, and decision variables of the execution are recorded and stored after running is completed. Based on these data, metric results, significance tests, and schematic plotting would be able to be continued. Metrics functions running on the data are divided into single-task and multitask metrics. Selecting single-task metrics displays the results of each task of the MTO problem on the table in MToP, and selecting multitask metrics displays the results of each MTO problem. The results of the significance tests are also shown in the table along with the metric results. All table data can be exported in tex, csv, and xlsx types as shown in Table V. Convergence plots are displayed based on the selected table data to analyze the comparison of the convergence behavior of different algorithms as in Fig. 3 (a). Additionally, for the multi-objective MTO algorithms and problems, the Pareto front can plotted as in Fig. 3 (b).
@@ II-A3 Data Process Module
TABLE VI: Properties of experimental data.
@@
Each unique execution in the experiment module is treated as a data object and saved to a data file MTOData.mat. The settings and intermediate results of this experiment are stored in data properties as shown in Table VI. The saved data can subsequently be reloaded by the experiment module of MToP. With the data process module, the data can be merged or split which allows data to be reused and customized. The processing of data is grouped into three types, which are independent runs, algorithms and problems. Having several sets of data with fewer independent runs, the data can then be merged to become new data with a larger number of independent runs. In cases where existing data needs to be augmented with new algorithms for comparison, with the data process module there is no need to re-execute all the comparative experiments. It is only necessary to run the new algorithms and merge the resulting data with the original data. Similarly, if not all algorithms in the data are needed, it can be split and re-merged. As with the data processing for algorithms, the data processing for problems is applied in cases where new problems need to be added or deleted.

### 09 Apr 2024 v2

@@ Abstract
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 10 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
@@ I Introduction
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6]. In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking [7]. A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
@@
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. EMT has found successful applications in various domains, including engineering scheduling [5, 8], nonlinear equation systems [9], feature selection in machine learning [10], anomaly detection [11], point cloud registration [12, 13], and reinforcement learning [14].
@@
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques [15, 16]. Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources [17, 18]. The first attempt of EMT can be traced back to the multifactorial EA [19], which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks [20, 21], or a multi-population framework allocating separate populations for each task [22, 23]. Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space [19, 21], autoencoding [24, 25], affine transformation [26, 27, 28], and adversarial generative models [29] have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment [20, 30, 31], knowledge selection [23, 32], and historical feedback [33, 34, 35], have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs. Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field. Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field. In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization. For instance, PlatEMO [36] is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems. EvoX [37] is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training. Evo-Panel [38] created a comprehensive dynamic visualization tool to illustrate the detailed procedures of different optimization algorithms in solving numerical benchmark functions. EDOLAB [39] is an optimization tool specifically for dynamic optimization.
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. These distinctive requirements pose unsolvable challenges for MTEA implementation in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45]. Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
@@
To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. The main contributions of MToP are summarized as follows:
@@
    MToP features a user-friendly graphical user interface (GUI) comprising test, experiment, and data-process modules. These modules facilitate researchers in understanding problem characteristics, conducting comparative experiments, solving problems in parallel, statistically analyzing results, plotting result figures, and managing experimental data. Moreover, MToP offers a modular implementation of algorithms, problems and performance metrics. On top of these, it provides an extensive public application programming interface (API) with template functions for population initialization, function evaluation, evolutionary operators, and environmental selection.
@@
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types (see Table IV). Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types (see Table V). In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications (see Table VII). Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities (see Table VIII).
@@
    MToP is designed to be easily extended, allowing for the seamless addition of new algorithms, problems, and metrics. By adhering to established coding patterns and implementing functionality based on the public API, new code can be seamlessly integrated and utilized within MToP. Given its status as a completely open-source project, researchers have the opportunity to leverage existing algorithms, problems, and metrics as the foundation for novel ideas. Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research [^1] MToP has been continuously updated since Sep. 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
The rest of this paper is organized as follows. Section II elaborates on the architecture of MToP. Section III describes approaches for performing experiments with MToP. Section IV introduces how to extend the contents of MToP. Finally, the conclusion and future work are given in Section V.
@@ II-A1 Test Module
(a) Function landscapes (1D)
(b) Function landscapes (2D)
(c) Feasible regions
(d) Pareto front
(e) Metric convergence
@@
The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms. It facilitates running the selected algorithm on the chosen problem within MToP. The module includes visualization functions for both single-objective and multi-objective MTO problems. In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig. 1 (a)-(b). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig. 1 (c). Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig. 1 (d). Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig. 1 (e).
(a) Gen=1
(b) Gen=100
(c) Gen=200
(d) Gen=500
(e) Gen=700
(f) Gen=1000
Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space with the number of generations.
(a) Gen=1
(b) Gen=100
(c) Gen=200
(d) Gen=500
(e) Gen=700
(f) Gen=1000
Figure 3: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the decision space with the number of generations.
In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior. Figures 2 and 3 illustrate an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space. Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=100,200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500,700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task. However, these two tasks do not precisely overlap in the decision space. This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
@@
The workflow of API calls for executing an algorithm on a problem is depicted in Fig. 4. The process initiates with the Start_Button event in MTO_GUI.m, triggering the parameter setting API Set_Parameter within the selected algorithm MFEA.m. Subsequently, it proceeds to the Run function of the algorithm, which initially invokes the Initialization API of the algorithm base class Algorithm.m, and then iterates through the function evaluation API Evaluation. During the initialization phase, the parameter settings of the problem CEC17_MTP1.m to be solved are obtained via the problem base class Problem.m. Similarly, the Evaluation function retrieves the specific optimization task object through the problem base class. Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function Obj.m via the Call_Metric API.
All functions within MToP that execute algorithms, including subsequent experiment modules, adhere to this procedure. Additionally, both algorithm and problem parameters can be directly modified through the GUI of MToP.
@@ II-A2 Experiment Module
TABLE I: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
@@
The experiment module offers functionalities for conducting comprehensive experiments and analyzing results. Users can execute multiple independent runs of various algorithms to tackle multiple MTO problems. Upon completion of the runs, data including experimental settings, objective values, and decision variables are recorded and stored. These data serve as the basis for generating metric results, conducting significance tests, and creating schematic plots.
Metrics functions operating on the data are categorized into single-task and multitask metrics. The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem. Significance test results are also included in the table alongside the metric results. All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table I.
Convergence plots are generated based on the selected table data to facilitate the comparison of convergence behaviors among different algorithms, as depicted in Fig. 5 (a). The convergence plot function can be applied to both single- and multi-objective problems. Furthermore, for multi-objective MTO algorithms and problems, the Pareto front can be plotted, as illustrated in Fig. 5 (b).
@@ II-A3 Data Process Module
TABLE II: Properties of experimental data.
@@
Each unique execution within the experiment module is considered a data object and is saved to a data file named MTOData.mat. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II. This saved data can subsequently be reloaded by the experiment module of MToP.
With the data process module, the data can be merged or split, enabling data reuse and customization. The processing of data is categorized into three types: independent runs, algorithms, and problems. If several sets of data with fewer independent runs are available, the data can be merged to create new data with a larger number of independent runs. Similarly, in cases where existing data needs augmentation with new algorithms for comparison, it is unnecessary to re-execute all comparative experiments. Instead, only the new algorithms need to be executed, and the resulting data can be merged with the original data. Likewise, if not all algorithms within the data are required, the data can be split and re-merged accordingly. The data processing for problems follows a similar approach, wherein new problems can be added or deleted as needed.